The Investment department of the Zambia Development Agency randomly selected 100 SMEs and examined them for profitability on their investment. The following results were obtained:

Number of Profits


1

2

3

Number of SMEs

36

40

19

5

Test at 5% significance level whether the distribution of profits follows a Poison distribution.


1
Expert's answer
2021-03-16T18:17:29-0400

Denote X as number of SMEs.

Estimation of parameter "\lambda" of Poisson distribution:

"\lambda = EX = \frac{1}{n}\sum_{i=0}^{\infty}\nu_iX_i = \frac{36*0 + 40 * 1 + 19 * 2 + 5 * 3}{100} = 0.93"

Distribution function:

"p_k = Pr(X = k) = e^{-\lambda} \frac{\lambda^k}{k!}"

Let's consider Pearson's test for parametric distribution with null hypothesis

"H_0" : X has poisson distribution.

Statistics:

"T = \sum_{i=0}^3\frac{(\nu_i - p_i n)^2}{np_i} = \frac{36 - e^{-0.93} * 100}{e^{-0.93} * 100} + ... = 0.836"

T should be distributed as "\Chi^2(k - 1 - l) = \Chi^2(4 - 1 - 1) = \Chi^2(2)" , where "l" is parameters count.

At 5% significance level for this distribution we have critical value 5.991 > 0.836.

Hence, at 5% significance level the distribution of profits follows a Poison distribution.